Please justify answer, this is for a mathematical proofs class.2. Determine the number of ways one can pick two cards from a standard 52-card deck, so that the first card is a spade and th
Answer #1

Solution terc we haue to cletermine, the number of Pick tuoo Cards feom stundard the first card is a Spule ways one Can 52 Ca
Answer #2

ANSWER :


N(First card spade excluding spade ace) = 12 

N(First card spade ace) = 1


N(Second card not ace | First card spade excluding spade ace) = 51 - 4 = 47

N(Second not ace | First card is spade ace) = 51 - 3 = 48 


So, N(First card spade and second card not ace)


= N(First card spade excluding spade ace) * N(Second card not ace | First card spade excluding spade ace)

+ N(First card spade ace) * N(Second not ace | First card is spade ace)

= 12 * 47 + 1 * 48 


= 612 (ANSWER).

answered by: Tulsiram Garg